# Quantum information scrambling in adiabatically-driven critical systems

### v1

@@ 3 Lipkin-Meshkov-Glick model
where $\hat{S}_{\alpha}=\sum_{k=1}^{N}\hat{\sigma}_{\alpha}/2$ are the collective spin-$1/2$ operators along $\alpha$-direction ($\alpha=x,y,z$). Since the $\hat{H}(g)$ commutes with $\hat{S}^{2}=\hat{S}_{x}^{2}+\hat{S}_{y}^{2}+\hat{S}_{z}^{2}$, we restrict ourselves to the sector of maximum spin, i.e. $J=N/2$, and work in the Dicke basis, $\hat{S}_{z}\left|J,M\rangle\right.=M\left|J,M\rangle\right.$ with $-J\leq M\leq J$, and $\hat{S}^{2}\left|J,M\rangle\right.=J(J+1)\left|J,M\rangle\right.$. For simplicity, we will consider $N$ even. In the $N\rightarrow\infty$, the previous Hamiltonian features a QPT at the critical value $g_{c}=1$ [48, 47], which divides the spectrum into two phases (see Fig. 1 for a representation of the energy spectrum for $N=300$ spins). For $g<g_{c}$, there is a symmetry-breaking $\mathbb{Z}_{2}$ phase up to critical energy $E_{c}(g)$ dependent on $g$ where the eigenstates of different parity are degenerate; to the contrary, for energies above $E_{c}(g)$ or for $g>g_{c}$, there is a normal phase where the parity is well defined and the eigenstates are no longer degenerate. The location of the critical excitation energy $E_{c}(g)$ corresponds to the so-called excited-state quantum phase transition (ESQPT), whose main hallmark consists in a logarithmically-diverging density of states at $E_{c}(g)$ [45, 36, 29, 77, 30, 31, 52, 78, 32, 33]. In the LMG model, the spin operator $\hat{S}_{x}$ plays the role of the operator $\hat{\mathcal{O}}$ introduced in Sec. 2, whose ground-state expectation value serves as a good order parameter for the ground-state QPT.
Figure 1: Sketch of protocol that leads to an adiabatic scrambling of information. Information is encoded in an initially symmetry-breaking state $\lvert\psi_{0}\rangle$, which populates a large collection of eigenstates in the symmetry-broken phase of the model at $g_{0}<g_{c}$. The system then undergoes a coherent adiabatic cycle $g(t)$ entering into the normal phase and back. The populations of the initial state remain unaltered, while the information is scrambled due to the change in the relative phases among degenerate eigenstates. The energy spectrum corresponds to the LMG for $N=300$ spins, where $\Delta_{0,k}(g)=E_{k}(g)-E_{0}(g)$ with dashed- and solid-lines depicting the energies of even and odd parity eigenstates, respectively.
In the following, we illustrate how the mechanism described in Sec. 2 applies to the LMG model, i.e. the adiabatic quantum information scrambling of the symmetry-breaking initial state. To study the effectiveness of AQIS as a function of the number of populated eigenstates, i.e. the validity of Eq. (10), we first consider an initial microcanonical-like state that uniformly populates the first $2N_{\rm mc}$ eigenstates. That is,
@@
where the coefficients $x_{k}=-1,+1$ are chosen according to the sign of $\langle\phi_{k,+}|\hat{S}_{x}|\phi_{k,-}\rangle$ so that $\left.\langle\psi_{0}\right|\hat{S}_{x}\left|\psi_{0}\rangle\right.$ is positive, and the probability distribution $P(S_{x})$ only shows non-zero values in the positive branch. See Fig. 2(a) and (b) for the initial probability distribution of this microcanonical-like state where $g_{0}=0$, $N=100$ spins with $N_{\rm mc}=10$. The dynamical protocol follows Eq. (5) with $g_{1}=1.25$ to ensure that the system is completely driven into the normal phase (cf. Fig. 1). The final probability distributions are also plotted in Fig. 2(a) and (b), for the $\hat{S}_{x}$ and energy, respectively. Considering slow ramps ($\tau\gg 1$) we first note that the energy probability distribution $P(E)$ remains unaltered, i.e. the coefficients $c_{k,p}$ approximately hold constant. To the contrary, $P(S_{x})$ significantly differs from the initial distribution as a consequence of the AQIS. The final state $\left|\psi(2\tau)\rangle\right.$ populates both branches of the observable $\hat{S}_{x}$, and thus its expectation value after the completion of the cycle vanishes, $\left.\langle\psi(2\tau)\right|\hat{S}_{x}\left|\psi(2\tau)\rangle\right.\approx 0$. This can be better visualized in Fig. 2(c) where we show the evolution of $\left.\langle\psi(t)\right|\hat{S}_{x}\left|\psi(t)\rangle\right.$ as a function of $g(t)$. Since the dynamical evolution is to a very good approximation adiabatic, other observables not related to the symmetry-breaking information remain unchanged, such as $\hat{S}_{z}$ or the energy (cf. Fig. 2(d)). In the following, and to ease the numerical simulations, we assume that the evolution is adiabatic and thus make use of the approximation given in Eq. (6) to compute the final state after the cycle.
Figure 2: Adiabatic quantum scrambling of the symmetry-breaking information in the LMG model. Panel (a) shows the initial and final probability distribution over the observable $\hat{S}_{x}$ for an initial microcanonical-like state undergoing a fully-coherent cycle from the symmetry-breaking to the normal phase and back. Panel (b) shows the energy probability distribution $P(E)$ for the initial and final states after the cycle (same $\tau$ values as in (a)). Panel (c) illustrates the dynamics of the expectation value $\left.\langle\psi(t)\right|\hat{S}_{x}\left|\psi(t)\rangle\right.$ as a function of the time-dependent control $g(t)$ for $\tau=500$. Here $F$ ($B$) denotes the forward (backward) process from $g_{0}\rightarrow g_{1}$ ($g_{1}\rightarrow g_{0}$). Note that $\langle S_{x}\rangle\approx 0$ after the adiabatic protocol. In (d) the dynamics of other relevant observables not related to the symmetry-breaking information are shown, and thus not scrambled during the adiabatic protocol. The employed parameters are $N=100$ spins, $g_{0}=0$, $g_{1}=1.25$, $N_{\rm mc}=10$.
@@ 3.1 Effectiveness of the quantum information scrambling
The effectiveness of the AQIS of the symmetry-breaking information is intimately related to the distribution of the phases $\delta\varphi_{k}$, i.e. the difference of the accumulated phase among eigenstates of opposite parity, as discussed in Sec. 2. Thus, the proper quantum information scrambling requires $\delta\varphi_{k}$ to uniformly sample $[0,2\pi)$. In Fig. 3 we show the resulting distribution of these phases $\delta\varphi_{k}$ for a LMG model comprising $N=2000$ spins and for the first 200 eigenstates and $\tau=10^{3}$ with $g_{0}=0$ and $g_{1}=1.25$. Note that it approximately corresponds to a uniform distribution in the range $[0,2\pi)$. Similar results can be found for other choices of $\tau$, $N$, $g_{0}$, and $g_{1}$, provided $g_{1}$ ensures that the system enters in the normal phase. Otherwise, i.e. if the system remains in the symmetry-breaking phase during the whole cycle, then $\delta\varphi_{k}=0\ \forall k$ since $E_{k,+}(g)=E_{k,-}(g)$ during the protocol and the information scrambling will not occur.
@@
As commented in Sec. 2, the effectiveness of AQIS depends on the number of populated eigenstates in the initial state, i.e. on the support of $\left|\psi_{0}\rangle\right.$ on the eigenstates of $\hat{H}(g_{0})$. The initial state given in Eq. (12) allows us to systematically analyze the validity of Eq. (10) as $N_{\rm mc}$ increases. We first note that, if $N_{\rm mc}\approx 1$ the final expectation value of $\hat{S}_{x}$ will simply undergo an oscillatory dependent on $\tau$. This can be seen in Fig. 4(a) for an initial state with $N_{\rm mc}=4$, which hardly scrambles the symmetry-breaking information as one can recover the initial value $\left.\langle\psi_{0}\right|\hat{S}_{x}\left|\psi_{0}\rangle\right.$ at a suitable time $\tau$. Yet, as $N_{\rm mc}$ grows, this is no longer the case due to the uniform distribution of the phases $\delta\varphi_{k}$, which forces the final expectation value to average to zero, i.e. $\left.\langle\psi(2\tau)\right|\hat{S}_{x}\left|\psi(2\tau)\rangle\right.\approx 0$. The effectiveness of AQIS can be therefore quantified in terms of the fluctuations of $\left.\langle\psi(2\tau)\right|\hat{S}_{x}\left|\psi(2\tau)\rangle\right.$ around its average value, i.e.
@@
In the previous expression $\overline{X}$ corresponds to the average value of $X$ in an interval of quench times $\tau\in[\tau_{0},\tau_{1}]$, where $\tau_{0}$ already ensures the validity of the adiabatic approximation. If the information is perfectly scrambled, then $\sigma^{2}(S_{x})=0$ so that the adiabatically-evolved final state features $\left.\langle\psi(2\tau)\right|\hat{S}_{x}\left|\psi(2\tau)\rangle\right.=0$ independently of $\tau$. Therefore, the smaller $\sigma(S_{x})$ the more effective is the AQIS. The results are plotted in Fig. 4(b) obtained for $\tau_{0}=10^{3}$ and $\tau_{1}=10^{4}$, and similar parameters as in previous figures, namely, $g_{0}=0$, $g_{1}=1.25$ and for two system sizes, $N=1000$ and $2000$ spins. As the number of populated states grows, $N_{\rm mc}$, the variance $\sigma^{2}(S_{x})$ decreases, thus indicating a better-scrambling performance. A fit to the numerical results reveals a dependence $\sigma(S_{x})\propto N_{\rm mc}^{-2/3}$ (cf. Fig. 4(b)).
Figure 4: Effectiveness of the AQIS as a function of the number of populated eigenstates. Panel (a) shows the final expectation value of $\hat{S}_{x}$, i.e. $\left.\langle\psi(2\tau)\right|\hat{S}_{x}\left|\psi(2\tau)\rangle\right.$ as a function of $\tau$, ensuring the adiabatic approximation, and for a different number of populated eigenstates $N_{\rm mc}$, i.e. $4$, $64$ and $512$. Panel (b) shows the the resulting standard deviation of the results in (a) for their average value, as a function of $N_{\rm mc}$. See the main text for further details.
@@ 3.2 Loschmidt echo and out-of-time-ordered correlator
which quantifies the overlap between the final state $\left|\psi(2\tau)\rangle\right.=\hat{U}(2\tau)\left|\psi_{0}\rangle\right.$ and the potentially restored initial state upon a time-reversal evolution with a small time-delay $\delta t$. Hence, for $\delta t=0$ it follows $L(\tau,\delta t=0)=1\ \forall\tau$. The scrambling due to AQIS will be robust the smaller the Loschmidt echo for small time-delays, i.e. if $L(\tau,\delta t)\approx 0$ for $\delta t\ll 1$.
Figure 5: Loschmidt echo and OTOC for the LMG model. Panels (a), (b), and (c) show the Loschmidt echo $L(\delta t)$ as a function of the time mismatch $\delta t$ and for initial states with different populated eigenstates $N_{\rm mc}$. Note that strong revivals even at late times (b), and the decaying echo $1/\delta t$ (dashed line) in (c) when $N_{\rm mc}\gg 1$. Panel (d) corresponds to a rescaled OTOC which signals a similar behavior to the expectation value of $S_{x}$. The simulation parameters are $N=2000$ spins, $g_{0}=0$ and $g_{1}=1.25$. See the main text for further details.
As we are interested in an adiabatic evolution, i.e. for sufficiently long $\tau$, we rely again on the adiabatic approximation. It is easy to see that under this approximation one obtains the simple expression
@@
with $\sum_{k,p}|c_{k,p}|^{2}=1$. Note that under the adiabatic approximation, the previous quantity does not depend on $\tau$, i.e. $L(\tau,\delta t)\equiv L(\delta t)$. Again, as for the effectiveness of AQIS (Eq. (10)), initial states with a large support and uniformly distributed phases $\varphi_{k,p}$, one expects $L(\delta t)\approx 0$ even for $0<\delta t\ll 1$. The results are plotted in Fig. 5(a)-(c), which precisely reveal a vanishingly small echo $L(\delta t)\approx 0$ for small time mismatches $\delta t$. Although not explicitly shown, we note that similar results can be obtained if instead of a time mismatch $\delta t$ a deviation in other quantity is considered, such as in $g_{0,1}$. In particular, we observe a decay $L(\delta t)\propto 1/\delta t$ for short $\delta t$ and $N_{\rm mc}\gg 1$. As expected, in states with small support (few populated eigenstates), $L(\delta t)$ displays large revivals and never decays to 00, a clear indication of the failure of quantum information scrambling.
As commented in Sec. 1, OTOCs have been proven a valuable tool for studying information scrambling. In a sudden quench scenario, scrambling is related to the short-time behavior of the OTOC for a relevant observable [26]. In our case, we can define the OTOC as
@@
Since the AQIS is effective only after the cycle has been completed, we focus on $O(2\tau)$. Again, relying on the adiabatic approximation of the evolved state, we approximate $\hat{U}(2\tau)$ by its adiabatic counterpart $\hat{U}^{\rm Ad}(2\tau)$, so the populations in the eigenbasis of the instantaneous Hamiltonian remain constant and the only change is due to the accumulated phase during the cycle. In this manner, we can compute the adiabatic OTOC, denoted here as $O^{\rm Ad}(2\tau)$ and fulfilling $O^{\rm Ad}(2\tau)\approx O(2\tau)$ for $\tau\gg 1$. The results for $O^{\rm Ad}(2\tau)$, plotted in Fig. 5(d), reveal essentially the same behavior as $\left.\langle\psi(2\tau)\right|\hat{S}_{x}\left|\psi(2\tau)\rangle\right.$. That is, the adiabatic dynamics scramble the symmetry-breaking information contained in the initial state, and thus $O(2\tau)\approx 0$. This becomes more effective the more eigenstates are significantly populated in the initial state. Conversely, for small support (e.g. $N_{\rm mc}=4$ corresponding to the red line in Fig. 5(d)), the OTOC features large oscillations, potentially reaching its initial value.
Figure 6: AQIS for a symmetry-breaking thermal state in the LMG model. (a) Initial probability distribution $P(S_{x})$ over the observable $\hat{S}_{x}$ for $\hat{\rho}_{\beta,SB}$ with $\beta=0.02$ for $N=10^{3}$ and $g_{0}=0$ (solid black line), together with the probability for the state upon an adiabatic protocol for $\tau=10^{3}$ and $g_{1}=1.25$ (blue squares), which equally populates both $S_{x}$ branches. (b) Population of the initial eigenstate for the symmetry-breaking thermal state $\hat{\rho}_{\beta,SB}$, as in (a). (c) Final expectation value of $\hat{S}_{x}$ after an adiabatic protocol from $g_{0}=0$ to $g_{1}=0.1$ (red), $0.65$ (green) and $1.25$ (blue), as a function of $\tau$ and for $\hat{\rho}_{\beta,SB}$ with $\beta=0.02$ and $N=10^{3}$, as in (a) and (b). For $g_{1}=0.1$, the system is not driven into the normal phase. For $g=0.65$, the state enters partially, while for $g_{1}=1.25$ it is driven completely into the normal phase, leading to a full AQIS. (d) Average expectation value of $\hat{S}_{x}$ in the range $\tau\in\{10^{3},10^{4}\}$ as a function of the final coupling strength $g_{1}$, revealing an order-parameter-like behavior. The error bars correspond to a standard deviation for the average of $\langle\hat{S}_{x}(2\tau)\rangle$. See main text for details.

### v2

@@ 3 Lipkin–Meshkov–Glick Model
where $\hat{S}_{\alpha}=\sum_{k=1}^{N}\hat{\sigma}_{\alpha}/2$ are the collective spin-$1/2$ operators along the $\alpha$-direction ($\alpha=x,y,z$). Since the $\hat{H}(g)$ commutes with $\hat{S}^{2}=\hat{S}_{x}^{2}+\hat{S}_{y}^{2}+\hat{S}_{z}^{2}$, we restrict ourselves to the sector of maximum spin, i.e., $J=N/2$, and work in the Dicke basis, $\hat{S}_{z}\left|J,M\rangle\right.=M\left|J,M\rangle\right.$ with $-J\leq M\leq J$, and $\hat{S}^{2}\left|J,M\rangle\right.=J(J+1)\left|J,M\rangle\right.$. For simplicity, we will consider $N$ to be even. In the $N\rightarrow\infty$, the previous Hamiltonian features QPT at the critical value $g_{c}=1$ [60, 59], which divides the spectrum into two phases (see Figure 1 for a representation of the energy spectrum for $N=300$ spins). For $g<g_{c}$, there is a symmetry-breaking $\mathbb{Z}_{2}$ phase up to critical energy $E_{c}(g)$, dependent on $g$, where the eigenstates of different parity are degenerate; to the contrary, for energies above $E_{c}(g)$ or for $g>g_{c}$, there is a normal phase where the parity is well defined and the eigenstates are no longer degenerate. The location of the critical excitation energy $E_{c}(g)$ corresponds to the so-called excited-state quantum phase transition (ESQPT), whose main hallmark consists of a logarithmically diverging density of states at $E_{c}(g)$ [57, 48, 41, 88, 42, 43, 64, 89, 44, 45]. In the LMG model, the spin operator $\hat{S}_{x}$ plays the role of the operator $\hat{\mathcal{O}}$ introduced in Section 2, whose ground-state expectation value serves as a good order parameter for the ground-state QPT.
In the following, we illustrate how the mechanism described in Section 2 applies to the LMG model, i.e., the adiabatic quantum information scrambling of the symmetry-breaking initial state. To study the effectiveness of AQIS as a function of the number of populated eigenstates, i.e., the validity of Equation (10), we first consider an initial microcanonical-like state that uniformly populates the first $2N_{\rm mc}$ eigenstates. That is,
@@
where the coefficients $x_{k}=-1,+1$ are chosen according to the sign of $\langle\phi_{k,+}|\hat{S}_{x}|\phi_{k,-}\rangle$, so that $\left.\langle\psi_{0}\right|\hat{S}_{x}\left|\psi_{0}\rangle\right.$ is positive, and the probability distribution $P(S_{x})$ only shows non-zero values in the positive branch. See Figure 2a,b for the initial probability distribution of this microcanonical-like state where $g_{0}=0$ and $N=100$ spins with $N_{\rm mc}=10$. The dynamical protocol follows Equation (5) with $g_{1}=1.25$ to ensure that the system is completely driven into the normal phase (cf. Figure 1). The final probability distributions are also plotted in Figure 2a,b for the $\hat{S}_{x}$ and energy, respectively. Considering slow ramps ($\tau\gg 1$), we first note that the energy probability distribution $P(E)$ remains unaltered, i.e., the coefficients $c_{k,p}$ approximately hold constant. To the contrary, $P(S_{x})$ significantly differs from the initial distribution as a consequence of AQIS. The final state $\left|\psi(2\tau)\rangle\right.$ populates both branches of the observable $\hat{S}_{x}$, and thus its expectation value after the completion of the cycle vanishes, $\left.\langle\psi(2\tau)\right|\hat{S}_{x}\left|\psi(2\tau)\rangle\right.\approx 0$. This can be better visualized in Figure 2c, where we show the evolution of $\left.\langle\psi(t)\right|\hat{S}_{x}\left|\psi(t)\rangle\right.$ as a function of $g(t)$. Since the dynamical evolution is to a very good approximation adiabatic, other observables not related to the symmetry-breaking information remain unchanged, such as $\hat{S}_{z}$ or the energy (cf. Figure 2d). In the following, and to ease the numerical simulations, we assume that the evolution is adiabatic, and thus make use of the approximation given in Equation (6) to compute the final state after the cycle.
Figure 1: Sketch of protocol that leads to an adiabatic scrambling of information. Information is encoded in an initially symmetry-breaking state $\lvert\psi_{0}\rangle$, which populates a large collection of eigenstates in the symmetry-broken phase of the model at $g_{0}<g_{c}$. The system then undergoes a coherent adiabatic cycle $g(t)$, entering into the normal phase and back. The populations of the initial state remain unaltered, while the information is scrambled due to the change in the relative phases among degenerate eigenstates. The energy spectrum corresponds to the LMG for $N=300$ spins, where $\Delta_{0,k}(g)=E_{k}(g)-E_{0}(g)$, with the dashed and solid lines depicting the energies of even and odd parity eigenstates, respectively.
@@ 3.1 Effectiveness of the Quantum Information Scrambling
The effectiveness of the AQIS of the symmetry-breaking information is intimately related to the distribution of the phases $\delta\varphi_{k}$, i.e., the difference of the accumulated phase among eigenstates of opposite parity, as discussed in Section 2. Thus, the proper quantum information scrambling requires $\delta\varphi_{k}$ to uniformly sample $[0,2\pi)$. In Figure 3, we show the resulting distribution of these phases $\delta\varphi_{k}$ for a LMG model comprising $N=2000$ spins for the first 200 eigenstates, and $\tau=10^{3}$ with $g_{0}=0$ and $g_{1}=1.25$. Note that it approximately corresponds to a uniform distribution in the range $[0,2\pi)$. Similar results can be found for other choices of $\tau$, $N$, $g_{0}$, and $g_{1}$, provided that $g_{1}$ ensures that the system enters in the normal phase. Otherwise, if the system remains in the symmetry-breaking phase during the whole cycle, then $\delta\varphi_{k}=0\ \forall k$, since $E_{k,+}(g)=E_{k,-}(g)$ during the protocol, and information scrambling will not occur.
Figure 2: Adiabatic quantum scrambling of the symmetry-breaking information in the LMG model. Panel (a) shows the initial and final probability distribution over the observable $\hat{S}_{x}$ for an initial microcanonical-like state undergoing a fully coherent cycle from the symmetry-breaking to the normal phase and back. Panel (b) shows the energy probability distribution $P(E)$ for the initial and final states after the cycle (same $\tau$ values as in (a)). Panel (c) illustrates the dynamics of the expectation value $\left.\langle\psi(t)\right|\hat{S}_{x}\left|\psi(t)\rangle\right.$ as a function of the time-dependent control $g(t)$ for $\tau=500$. Here, $F$ ($B$) denotes the forward (backward) process from $g_{0}\rightarrow g_{1}$ ($g_{1}\rightarrow g_{0}$). Note that $\langle S_{x}\rangle\approx 0$ after the adiabatic protocol. In (d) the dynamics of other relevant observables not related to the symmetry-breaking information are shown, and thus not scrambled during the adiabatic protocol. The employed parameters are $N=100$ spins, $g_{0}=0$, $g_{1}=1.25$, $N_{\rm mc}=10$.
@@
As stated in Section 2, the effectiveness of AQIS depends on the number of populated eigenstates in the initial state, i.e., on the support of $\left|\psi_{0}\rangle\right.$ on the eigenstates of $\hat{H}(g_{0})$. The initial state given in Equation (12) allows us to systematically analyze the validity of Equation (10) as $N_{\rm mc}$ increases. We first note that, if $N_{\rm mc}\approx 1$, then the final expectation value of $\hat{S}_{x}$ will simply undergo an oscillatory dependent on $\tau$. This can be seen in Figure 4a for an initial state with $N_{\rm mc}=4$, which hardly scrambles the symmetry-breaking information, as one can recover the initial value $\left.\langle\psi_{0}\right|\hat{S}_{x}\left|\psi_{0}\rangle\right.$ at a suitable time $\tau$. Yet, as $N_{\rm mc}$ grows, this is no longer the case, due to the uniform distribution of the phases $\delta\varphi_{k}$, which forces the final expectation value to average to zero, i.e., $\left.\langle\psi(2\tau)\right|\hat{S}_{x}\left|\psi(2\tau)\rangle\right.\approx 0$. The effectiveness of AQIS can therefore be quantified in terms of the fluctuations of $\left.\langle\psi(2\tau)\right|\hat{S}_{x}\left|\psi(2\tau)\rangle\right.$ around its average value, i.e.,
@@
In the previous expression, $\overline{X}$ corresponds to the average value of $X$ in an interval of quench times $\tau\in[\tau_{0},\tau_{1}]$, where $\tau_{0}$ already ensures the validity of the adiabatic approximation. If the information is perfectly scrambled, then $\sigma^{2}(S_{x})=0$, so that the adiabatically evolved final state features $\left.\langle\psi(2\tau)\right|\hat{S}_{x}\left|\psi(2\tau)\rangle\right.=0$ independently of $\tau$. Therefore, the smaller the $\sigma(S_{x})$, the more effective AQIS is. The results are plotted in Figure 4b, obtained for $\tau_{0}=10^{3}$, $\tau_{1}=10^{4}$, and similar parameters as in the previous figures, namely $g_{0}=0$, $g_{1}=1.25$, and for two system sizes, $N=1000$ and $2000$ spins. As the number of populated states grows, $N_{\rm mc}$, the variance $\sigma^{2}(S_{x})$ decreases, thus indicating a better scrambling performance. A fit to the numerical results reveals a dependence $\sigma(S_{x})\propto N_{\rm mc}^{-2/3}$ (cf. Figure 4b).
Figure 4: Effectiveness of AQIS as a function of the number of populated eigenstates. Panel (a) shows the final expectation value of $\hat{S}_{x}$, i.e., $\left.\langle\psi(2\tau)\right|\hat{S}_{x}\left|\psi(2\tau)\rangle\right.$ as a function of $\tau$, ensuring the adiabatic approximation, and for a different number of populated eigenstates $N_{\rm mc}$, i.e., $4$, $64$ and $512$. Panel (b) shows the the resulting standard deviation of the results in (a) for their average value, as a function of $N_{\rm mc}$. See the main text for further details.
@@ 3.2 Loschmidt Echo and Out-of-Time-Ordered Correlator
which quantifies the overlap between the final state $\left|\psi(2\tau)\rangle\right.=\hat{U}(2\tau)\left|\psi_{0}\rangle\right.$ and the potentially restored initial state upon a time-reversal evolution with a small time-delay $\delta t$. Hence, for $\delta t=0$, it follows that $L(\tau,\delta t=0)=1\ \forall\tau$. The scrambling due to AQIS will be robust the smaller the Loschmidt echo for small time-delays, i.e., if $L(\tau,\delta t)\approx 0$ for $\delta t\ll 1$.
As we are interested in an adiabatic evolution, i.e., for sufficiently long $\tau$, we rely again on the adiabatic approximation. It is easy to see that, under this approximation, one obtains the simple expression
@@
with $\sum_{k,p}|c_{k,p}|^{2}=1$. Note that under the adiabatic approximation, the previous quantity does not depend on $\tau$, i.e., $L(\tau,\delta t)\equiv L(\delta t)$. Again, as for the effectiveness of AQIS (Equation (10)), for initial states with a large support and uniformly distributed phases $\varphi_{k,p}$, one expects $L(\delta t)\approx 0$, even for $0<\delta t\ll 1$. The results are plotted in Figure 5a–c, which precisely reveal a vanishingly small echo $L(\delta t)\approx 0$ for small time mismatches $\delta t$. Although not explicitly shown, we note that similar results can be obtained if, instead of a time mismatch $\delta t$, a deviation in another quantity is considered, such as in $g_{0,1}$. In particular, we observe a decay $L(\delta t)\propto 1/\delta t$ for short $\delta t$ and $N_{\rm mc}\gg 1$. As expected, in states with little support (few populated eigenstates), $L(\delta t)$ displays large revivals, and never decays to 00, a clear indication of the failure of quantum information scrambling.
Figure 5: Loschmidt echo and OTOC for the LMG model. Panels (ac) show the Loschmidt echo $L(\delta t)$ as a function of the time mismatch $\delta t$ and for initial states with different populated eigenstates $N_{\rm mc}$. Note strong revivals even at late times (b), and the decaying echo $1/\delta t$ (dashed line) in (c) when $N_{\rm mc}\gg 1$. Panel (d) corresponds to a rescaled OTOC, which signals a similar behavior to the expectation value of $S_{x}$. The simulation parameters are $N=2000$ spins, $g_{0}=0$, and $g_{1}=1.25$. See the main text for further details.
As commented in Section 1, OTOCs have been proven a valuable tool for studying information scrambling. In a sudden quench scenario, scrambling is related to the short-time behavior of the OTOC for a relevant observable [29]. In our case, we can define the OTOC as
@@
Since AQIS is effective only after the cycle has been completed, we focus on $O(2\tau)$. Again, relying on the adiabatic approximation of the evolved state, we approximate $\hat{U}(2\tau)$ by its adiabatic counterpart $\hat{U}^{\rm Ad}(2\tau)$, so that the populations in the eigenbasis of the instantaneous Hamiltonian remain constant, and the only change is due to the accumulated phase during the cycle. In this manner, we can compute the adiabatic OTOC, denoted here as $O^{\rm Ad}(2\tau)$ and fulfilling $O^{\rm Ad}(2\tau)\approx O(2\tau)$ for $\tau\gg 1$. The results for $O^{\rm Ad}(2\tau)$, plotted in Figure 5d, reveal essentially the same behavior as $\left.\langle\psi(2\tau)\right|\hat{S}_{x}\left|\psi(2\tau)\rangle\right.$. That is, the adiabatic dynamics scramble the symmetry-breaking information contained in the initial state, and thus $O(2\tau)\approx 0$. This becomes more effective the more eigenstates are significantly populated in the initial state. Conversely, for small support (e.g., $N_{\rm mc}=4$ corresponding to the red line in Figure 5d), the OTOC features large oscillations, potentially reaching its initial value. Although not explicitly shown here, it is worth mentioning that the corresponding OTOC for operators not related to the symmetry-breaking information, such as $\hat{S}_{z}$, do not show scrambling. As a consequence, these OTOCs are approximately constant, and close to their initial value.